OpenStudy (anonymous):

HELP. Mr Brown asks a garden designer to make a square flower bed. he cost of paving the border of the flower bed is $20 per metre. The cost of the flowers and soil in the flower bed is $60 per m^2. If the total cost of making the flower bed is $3500, find the length of a side of the flower bed,

OpenStudy (anonymous):

Draw a square and call a side of it x. Then u have 2x worth of border and x^2 worth of bed and the sum of the two is 3,500. Solve the equation u get.

OpenStudy (anonymous):

:O omg thank u . i didn't realise that

OpenStudy (anonymous):

um im a little stucked. could u help me solve?

OpenStudy (anonymous):

OK, you have x^2 + 2x = 3500 which is a quadratic, you know the quadratic formula?

OpenStudy (anonymous):

The one that goes - b plus minus sqrt(.......

OpenStudy (anonymous):

If you don't know it, I will show u...

OpenStudy (anonymous):

okay i get it :D my answer is 7m

OpenStudy (anonymous):

Let's see 4x of border is 28*20 = 560 and 49*60 is 2940 for a total of 3500 looks good to me. (Even though I misled you about the border, putting 2x instead of 4x:-)

OpenStudy (anonymous):

:D haha. thanks

OpenStudy (anonymous):

Did you do it with quadratic or just sort of guess it?

OpenStudy (anonymous):

quadratic equation. :) heres my working. let 1 side of the square flower bed be x cost of 4 sides=$80x cost of area=$60x^2 60x^2+80x=3500 60x^2+80x-3500=0 (10x-70)(6x+50)=0 either 10x-70=0 or 6x+50=0 x=7 or x=-8 1/3 (rejected) my answer is 7m

OpenStudy (anonymous):

Perfect:-)

OpenStudy (anonymous):

thanks for the medal :)

OpenStudy (anonymous):

Deserved.

OpenStudy (anonymous):

:) hey can i ask u smth? when do we use quadratic to solve and when do we use simultaneous(elimination and substitution method)?

OpenStudy (anonymous):

The essential difference is the number of variables (unknowns). In the above we only had one, x. If you have two unknowns, then you need two equations (and three for three,etc) and then you would solve these by some method like elimination or substitution. In fact what you are doing is using the info you have to reduce the two equations in 2 unknowns to a single equation in 1 unknown which you can solve (and then substitute back to get the other unknown).

OpenStudy (anonymous):

okay :) so does that mean that we use simultaneous(substitution n elimination) for equations with more than 1 variable? & quadratic for equations with 1 variable?

OpenStudy (anonymous):

Roughly speaking, yes (of course some equations in 1 variable do not involve a squared term like x +3 = 4, linear as opposed to quadratic, or may involve higher powers, like a cubic). You will also learn some other methods when there is more than 1 variable.

OpenStudy (anonymous):

o.O yeahh okay :) thanks and see u!

OpenStudy (anonymous):

And you, ciao!